How do I work out this problem 1/2+(2/3 / 3/4)-(4/5*5/6)=5/6
Answer by Alan3354(69443)

How do I work out this problem 1/2+(2/3 / 3/4)-(4/5*5/6)=5/6
-------------
It appears to be worked out.
We can check it, tho.
1/2+(2/3 / 3/4)-(4/5*5/6)
= 1/2 + (2/3)*(4/3) - 20/30
= 1/2 + 8/9 - 2/3
= 1/2 + 8/9 - 6/9
= 1/2 + 2/9
= 9/18 + 4/18
= 13/18
------
It's not 5/6